The average molecular speed in a sample of CH4 gas at a certain temperature is 659 m/s.

The average molecular speed in a sample of H2 gas is ____________
m/s at the same temperature.
1
Expert's answer
2020-06-08T15:23:03-0400

The average molecular speed formula is

"v= \sqrt{\frac{8RT}{\pi M}}" where v is inversely proportional to M 1/2


Therefore,

"\dfrac{V1}{V2}" = "[\dfrac{M2}{M1}]^1/2"

"\dfrac{659}{V2}=" "[\dfrac{2}{16}]^1/2"


"V2=\dfrac{659}{0.015625}"


"V2=42176 m\/s"